OpenStudy (anonymous):

solve the absolute value equation. 4|4x-3|+3=-9

Directrix (directrix):

4|4x-3| + 3 = -9 Add (-3) to both sides of the equation and post what you get. Then, we will decide what to do next.

OpenStudy (anonymous):

Subtract 3 from both sides then divide both sides by 4 and then you can take it from there right?

Directrix (directrix):

@nicole-0109 What I thought you might post is what you get after you subtract 3 (which is the same as adding -3). Would you do that now? 4|4x-3| + 3 = -9 -3 -3 ------------------- ? ?

Directrix (directrix):

@nicole-0109 I can't just drop the answer on you without making an effort to help you understand the answer. If you don't want to work together, then I can't help you. Sorry.